OpenStudy (anonymous):

solve for x 8-3x-x=12 a.-1 b.-10 c.-2 d. -5

OpenStudy (anonymous):



OpenStudy (jadeishere):

8 - 3x - x = 12 -4x + 8 = 12 - 8 - 8 -4x = 12 - 8 -4x = 4 x = -1

OpenStudy (anonymous):

thx

OpenStudy (jadeishere):

Yeppers :)